For the transmission line per unit length parameters are 0.1 Ω/m, 0.2 μH/m, 10 pF/m and 0.02



℧/m. Find the complex propagation constant at a) 1 MHz b) 1 GHz



(Hint. γ = (𝑅 + 𝑗ω𝐿) × (𝐺 + 𝑗ω𝐶) )

1
Expert's answer
2022-01-13T09:22:08-0500

"\gamma=\sqrt{(R+j\omega L)(G+j\omega C)},"

"\gamma_1=0.7+0.3j,"

"\gamma_2=0.9+0.6j."